# Qubit error bursts in superconducting quantum processors of Quantum Inspire: quasiparticle pumping and anomalous time dependence

### v1

@@ Abstract
The increased propensity for qubits to relax, made global by phonon propagation, is a source of correlated errors and consequently problematic for quantum error correction (QEC) and eventually fault-tolerant quantum computing. In fact, such error bursts already set the performance limit of state-of-the-art QEC experiments [9, 10]. Several strategies are being pursued that can help mitigate the impact. Recent large-scale efforts involve moving experimental setups to deep underground facilities to reduce the flux of ionizing radiation of cosmogenic origin [11, 12]. Device-level innovations include phonon traps [13, 14, 15, 16], substrate etching to block phonon propagation [17, 18, 19], backside metallization to enhance phonon downconversion [20, 21, 22], QP traps using normal and low-gap superconductors [5, 23, 24, 25, 26], and band-gap engineering of JJ electrodes [27, 7, 28]. These innovations are often assessed not by their impact on the rate or recovery dynamics of qubit error bursts but on the average QP density in qubit JJ electrodes, extracted by measuring QP tunneling rates [29]. Some experiments observe exponential or power-law time dependence of QP tunneling rates after cooldown [30, 31, 32, 33], attributed to gradual thermalization of substrates, metal films, device packaging, and shielding.
@@
with $\Delta_{\mathrm{Bulk}}=180\penalty 10000\ \mu\mathrm{eV}$ and $\alpha=600\penalty 10000\ \mu\mathrm{eV}\cdot\mathrm{nm}$, we estimate a gap difference between the layers $\delta\Delta_{\mathrm{S\text{-}5}}\sim h\times 2.4\penalty 10000\ \mathrm{GHz}$. As $hf_{\mathrm{q}}>\delta\Delta_{\mathrm{S\text{-}5}}$ for all its transmons, we do not expect band-gap engineering effects [27, 7, 28] to play a role in $\mathrm{S\text{-}5}$.
@@ I Acknowledgements
We thank J. Baselmans, K. Karatsu, C. Andersen, B. Plourde, L. Kouwenhoven, G. Frossati, and their teams for helpful discussions; M. Moreira for implementing mid-circuit measurement in Quantum Inspire; J. Marques for initial experiments; S. van der Meer for assistance with dilution refrigerators; G. Calusine and W. Oliver for providing traveling-wave parametric amplifiers; the TU Delft Reactor Institute for the lead shield; the Kavli Nanolab Delft staff for cleanroom support; and V. Sinha for project management assistance. This research is funded by the Dutch National Growth Funds and Quantum Delta NL (KAT-1 StartImpulse and Phase 2), Intel Corporation, the Netherlands Organization for Scientific Research (NWA.1292.19.194), and the Dutch Ministry of Economic Affairs [Allowance for Top Consortia for Knowledge and Innovation (TKI)]. M.S. acknowledges funding from the Google PhD fellowship program.
@@ II Authors Contribution
G.R.D.C., M.S. and L.D.C conceptualized the experiment. G.R.D.C. and M.S. performed the experiment and data analysis. N.M. fabricated $\mathrm{S\text{-}5}$, A.K. and M.F. fabricated $\mathrm{S\text{-}7}$. M.B., N.H. and L.D.C. designed the two processors. B.S. developed monitoring software using the Quantum Inspire Software Development Kit. S.V.S. performed post-fabrication trimming of $\mathrm{S\text{-}7}$ readout and Purcell resonators. G.R.D.C., M.S. and L.D.C. wrote the manuscript with feedback from all the coauthors. L.D.C. supervised the project. G.R.D.C. and M.S. contributed equally to this work.

### v2

@@ Abstract
The increased propensity for qubits to relax, made global by phonon propagation, is a source of correlated errors and consequently problematic for quantum error correction (QEC) and eventually fault-tolerant quantum computing. In fact, such error bursts already set the performance limit of state-of-the-art QEC experiments [9, 10]. Several strategies are being pursued that can help mitigate the impact. Recent large-scale efforts involve moving experimental setups to deep underground facilities to reduce the flux of ionizing radiation of cosmogenic origin [11, 12]. Device-level innovations include phonon traps [13, 14, 15, 16], substrate etching to block phonon propagation [17, 18, 19], backside metallization to enhance phonon downconversion [20, 21, 22], QP traps using normal and low-gap superconductors [23, 24, 5, 25, 26], and band-gap engineering of JJ electrodes [27, 28, 29, 7, 30, 31]. These innovations are often assessed not by their impact on the rate or recovery dynamics of qubit error bursts but on the average QP density in qubit JJ electrodes, extracted by measuring QP tunneling rates [32]. Some experiments observe exponential or power-law time dependence of QP tunneling rates after cooldown [33, 34, 35, 36], attributed to gradual thermalization of substrates, metal films, device packaging, and shielding.
@@
with $\Delta_{\mathrm{Bulk}}=180\penalty 10000\ \mu\mathrm{eV}$ and $\alpha=600\penalty 10000\ \mu\mathrm{eV}\cdot\mathrm{nm}$, we estimate a gap difference between the layers $\delta\Delta_{\mathrm{S\text{-}5}}\sim h\times 2.4\penalty 10000\ \mathrm{GHz}$. As $hf_{\mathrm{q}}>\delta\Delta_{\mathrm{S\text{-}5}}$ for all its transmons, we do not expect band-gap engineering effects [27, 28, 29, 7, 30, 31] to play a role in $\mathrm{S\text{-}5}$.
@@ I Acknowledgements
We thank J. Baselmans, K. Karatsu, C. Andersen, B. Plourde, L. Kouwenhoven, G. Frossati, and their teams for helpful discussions; S. van der Meer and M. Sarsby for assistance with dilution refrigerators; G. Calusine and W. Oliver for providing traveling-wave parametric amplifiers; the TU Delft Reactor Institute for the lead shield; the Kavli Nanolab Delft staff for cleanroom support; and V. Sinha for project management assistance. This research is funded by the Dutch National Growth Funds and Quantum Delta NL (KAT-1 StartImpulse and Phase 2), the European Union Flagship on Quantum Technology (OpenSuperQplus100, No. 101113946), Intel Corporation, the Dutch Ministry of Economic Affairs [Allowance for Top Consortia for Knowledge and Innovation (TKI)], and the Netherlands Organization for Scientific Research (NWA.1292.19.194). M.S. acknowledges funding from the Google PhD fellowship program.
@@ II Authors Contribution
G.R.D.C., M.S., J.F.M., and L.D.C conceptualized the experiment. G.R.D.C. and M.S. performed the experiment and data analysis. J.F.M. performed initial experiments. N.M. fabricated $\mathrm{S\text{-}5}$, A.K. and M.F. fabricated $\mathrm{S\text{-}7}$. M.B., N.H. and L.D.C. designed the two processors. M.S.M. implemented mid-circuit measurement and B.S. developed monitoring software in Quantum Inspire. S.V.S. performed post-fabrication trimming of $\mathrm{S\text{-}7}$ resonators. G.R.D.C., M.S. and L.D.C. wrote the manuscript with feedback from coauthors. L.D.C. supervised the project. G.R.D.C. and M.S. contributed equally to this work.